OpenStudy (anonymous):

If f(x)=7-x^2 and g(x)=x-4 ... [f*g](x)=?

OpenStudy (anonymous):

=f[g(x)] =f(g(x)= 7-(g... ? Is that right so far?

OpenStudy (anonymous):

[f*g](x)= 7- (x-4)^2 solve it furhter

OpenStudy (anonymous):

7-x^2-16?

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

[f*g](x)= 7- (x-4)^2 = 7-(x^2+16-8x) do it further now

OpenStudy (anonymous):

where did 8x come from?

OpenStudy (anonymous):

how would i get rid of the parentheses?

OpenStudy (anonymous):

just multiply the negative sign with each term inside the parantheses

OpenStudy (anonymous):

So... 7+x^2-112+56x

OpenStudy (anonymous):

sorry: 7-x^2

OpenStudy (anonymous):

7-x^2-16+8x

OpenStudy (anonymous):

how did you get 16 and 8? do you not multiply 7 to those numbers?

OpenStudy (anonymous):

no i m not multiplying 7 to those numbers. i got 16 and 8 by the following formula \[(a-b)^{2} = a ^{2} +b ^{2}-2ab\]

OpenStudy (anonymous):

so if I had [g#f](x) i would start it by saying x-4(7-x^2) right?

OpenStudy (anonymous):

# is supposed to be *

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

welcome